A motorist applies the two forces shown to loosenHelp with #19 please
Answer #1

The unit vectors for the forces are the direction cosines. The position vector of the force F is

            rOF = -1.333k ft.

The magnitude of the moment due to F is IMOFI = ex .(rOF x F) = 1.230F ft lb

The magnitude of the moment due to –F is IM-OFI = ex .(r-OF x -F) = 1.230F ft lb

The total moment about x-axis is Mx = 1.230 i^ + 1.230 i^ = 2.46 i^

From which, for a total magnitude of 32 ft lb, the force to be applied is

           F = 32/2.46 = 13 lb